Grade 12Mechanics

With what speed must a ball be thrown down for it to bounce back 10m higher than its original level

Geetha
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student
 
if there is no loss of energy then you have to give extra kinetic energy to achieve extra 10 "m height.
law of conservation of energy
1/2 mu.u + mgh =mg( h + 10)
1/2mu.u +mgh=mgh+10mg
1/2 u.u =10g
u.u = 20g
u = 14
 
Regards
Arun (askIITians forum expert)